OpenStudy (anonymous):

Solve for the equation for the indicated variable 2x + Y = 8;y A. Y = 2x - 8 B. Y = 8 - 2x C. Y = 8 + 2x D. Y = 4 - x And 5y = 5x - 10;y A. Y = x - 2 B. Y = x + 2 C. Y = X - 10 -------- 5 D. Y = 5(5x -10) Could you also explain to me how you found the answer, so It'd be easier on me for future problems. Thanks for your help and time :)

OpenStudy (anonymous):

well first case : 2x + Y = 8 Y = 8 - 2x second: 5y = 5x - 10 y = (5x - 10) / 5 y = x - 2

OpenStudy (anonymous):

it was a pleasure to help you

OpenStudy (anonymous):

Thank you so much for your help :)